Waymire Ltd. manufactures two products and currently allocates overhcad based on direct labour hours. For the coming year, the company is considering changing to an activity based costing system by allocating overhead based on 3 activities. The estimated amount of overhead traceable to the 3 activities for the coming year are given below Activity Purchase orders Maintenance Machine Cost driver Cost #oforders $873069 347 Product I Product 2 599 19282 2198 $448164 18218 hours Inspection #of S704819 1736 inspections Total estimated direct labour hours for the coming year are 27302 hours for Product 1 and 80250 hours for Product 2 What is the difference between total overhead allocated to Product 2 using activity based costing versus traditional costing? Note: a negative number indicates less total overhead allocated using activity based costing. Select one a. $334684 b. $-848996 c. $-334684 d. $-662744
Answer #1
Activity Cost Total activity Activity rate X Activity for Product 2 Overhead allocated to Product 2
Purchase orders 873069 946 922.91 599 552821
Maintenance 448164 37500 11.95 19282 230440
Inspection 704819 3934 179.16 2198 393796
Total 2026052 1177056
Traditional costing:
Total overhead 2026052
Divide by Total labor hours 107552 =27302+80250
Overhead rate 18.84
X Product 2 hours 80250
Overhead allocated 1511740
Difference = 1177056-1511740 = $-334684
Option C is correct